Solve Problems Using Linear Systems
17. Fran and Winston have a combined income of $80 000. One quarter of Winston's income is the same as one-sixth of Fran's income. How much does each person earn?
Thankssssssssss

Answer by scott8148(6628)

"Fran and Winston have a combined income of $80 000" ___ F + W = 80000 ___ W = 80000-F

"One quarter of Winston's income is the same as one-sixth of Fran's income" ___ W/4 = F/6

substituting ___ (80000-F)/4 = F/6 ___ 20000 - (F/4) = F/6

adding F/4 ___ 20000 = 5F/12

multiplying by 12/5 ___ 48000 = F

substituting ___ W = 80000 - (48000) ___ W = 32000